Objects with masses of 225 kg and a 525 kg are separated by 0.500 m. (a) Find the net gravitational force exerted by these objects on a 56.0-kg object placed midway between them. magnitude direction toward the 525-kg mass (b) At what position (other than infinitely remote ones) can the 56.0-kg object be placed so as to experience a net force of zero? Did you accidentally divide or take the inverse in your calculation? m from the 525-kg mass The response you submitted has the wrong sign. N
Answer #1

Let the 2 original masses be m1, m2 separated by distance d, and the third mass be m3 at distance x from m2, and G be the gravitational constant.

The resultant force F towards m2 is:

F = G m2 m3 / x^2 - G m1 m3 / (d - x)^2

G m3 [ m2 / x^2 - m1 / (d - x)^2 ] ...(1)

(a)

Putting x = d / 2 in (1):

F = 4G m3 (m2 - m1) / d^2

= 4 * 6.673 × 10^(- 11) * 56.0(525 - 225) / 0.500^2

= 1.794 * 10^(- 5) N.

(b)

Equating F to 0 in (1):

m2 / x^2 = m1 / (d - x)^2

m1 x^2 = m2(d^2 + x^2 - 2dx)

(m2 - m1)x^2 - 2d m2 x + m2 d^2 = 0

x = (2d m2 +/- sqrt[ 4d^2 m2^2 - 4m2 d^2 (m2 - m1) ]) / [ 2(m2 - m1)

= d( m2 +/- sqrt[ m1 m2 ]) / (m2 - m1)

The negative sign gives the root between 0 and 0.500:

x = 0.500[ 525 - sqrt(225 * 525) ] / (525 - 225)

= 0.302 m